Some pleated draperies must be twice as wide as the window they cover. How many square yards of material are needed to cover this kind of window 2 ft. wide and 3 ft. long?
answer in square yards

Answer by Fombitz(32388)

Convert from feet to yards.
3 ft=1 yd
2 ft=2/3 yd
A=1(2/3)=2/3 sq.yd.